OpenStudy (anonymous):

Using the following measurements, find the length of the leg of this right triangle: leg = 5 hypotenuse = 13 12 13 144

jimthompson5910 (jim_thompson5910):

we have this right triangle 

jimthompson5910 (jim_thompson5910):

using the pythagorean theorem, we get a^2 + b^2 = c^2 5^2 + x^2 = 13^2 25 + x^2 = 169 x^2 = 169 - 25 x^2 = 144 x = sqrt(144) x = 12 So the missing side is 12 

OpenStudy (anonymous):

Thank u @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome